please I need your help


1.Suppose Ethiopian Electric Light and Power Corporation (EELPC) is a multi plantmonopolist having two plants, Tekeze plant (plant1) and Fincha plant (Plant2). Theoperating costs of the two plants are given as follows:Tekeze Plant: TC1 = 10 Q12and where Q1-Amount of electric power produced inTekezeFincha plant: TC2 = 20 Q22Q2–amount of electric power produced in FinchaEELPC estimates the demand for electric power by the following functionP= 700–5Q where P-is price (total in million birr) per Giga watt andQ–is the total amount of Giga watt sold and Q = Q1 + Q2Note that a Giga watt of electric power, whether it comes from Fincha or Tekeze plant worthequal price

a) What level of output (electric power) should EELPC produce and what price per Kilowattshould it charge to maximize its profit?

b) How much of the total output should be produced in each plant?

c) Suppose that recently the Tekeze plant is suffering from siltation problem



1
Expert's answer
2021-05-03T10:48:27-0400

a) The level of output that EELPC should produce is:

MR = MC1 = MC2,

MR = TR'(Q) = 700 - 10Q,

MC1 = TC1'(Q1) = 20Q1,

MC2 = 40Q2

700 - 10Q1 = 20Q1,

30Q1 = 700,

Q1 = 23.33 units.

700 - 10Q2 = 40Q2,

50Q2 = 700,

Q2 = 14 units.

P = 700 - 5×(23.33 + 14) = 513.33.

b) The total outputs produced in each plant are: Q1 = 23.33 units, Q2 = 14 units.